The line through A(-1, -2) and B(-5, y) is perpendicular to a line with slope 4. How would I solve for y?
Answer by Boreal(15235)

slope of the line through A and B is -1/4, the negative reciprocal of 4, since perpendicular lines have the product of their slopes equal to -1.
so the slope between A and B is (y+2)/(-5+1)=-(1/4)
(y+2)/-4=-1/4
multiply through by -4
y+2=1
y=-1